Abdul, while driving to school, computes the average speed for his trip to be 20 km h-1. On his return trip along the same route, there is less traffic and the average speed is 40 km h-1. What is the average speed for Abdul’s trip?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The distance while driving from Home to School = S
Assume time taken by Abdul to cover this distance = t1
Distance Abdul covers while driving from School to Home = S
Assume time taken by Abdul to cover this distance = t2
Average speed from home to school v1av = 20 km h-1
Average speed from school to home v2av = 30 km h-1
Also Time taken from Home to School t1 =S / v1av
Similarly, Time taken from School to Home t2 =S/v2av
Total distance from home to school and backward = 2 S
Total time taken from home to school and backward (T) = S/20+ S/30
Therefore, Average speed (Vav) for covering total distance (2S) = Total Dostance/Total Time
= 2S / (S/20 +S/30)
= 2S / [(30S+20S)/600]
= 1200S / 50S
= 24 kmh-1

Thanks